OpenStudy (anonymous):

May I please have some help? A medal and a fan will be given. :) The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n^2 + 2,500n - 4,000 Part A: What are the zeroes of the above function and what do they represent? Show your work. Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. Part C: What is the axis of symmetry of the function P(n)?

zepdrix (zepdrix):

\[\Large\rm P(n)=-250n^2 + 2500n - 4000\]So the find `zeros`, we set the P(n) equal to zero.\[\Large\rm 0=-250n^2 + 2500n - 4000\]This will be easier to work with if we have no fancy coefficient on the squared term. So let's factor a -250 out of each term:\[\Large\rm 0=-250(n^2-10n+16)\]Then divide the -250 away,\[\Large\rm 0=n^2-10n+16\]It looks like this will factor nicely! Confused by any of those steps?

OpenStudy (anonymous):

Hold on one second, let me look really quickly. :)

OpenStudy (anonymous):

Nope! I understand for sure! So, I find the GCF for the 16?

zepdrix (zepdrix):

Factors of 16 that add to -10. So we want two numbers that multiply to give us positive, but add to give us negative, So both factors should be negative. Can you find the right combination? :o

OpenStudy (anonymous):

Yes, I can hehehehehe...give me 3 seconds.

OpenStudy (anonymous):

\[-8 - 2 = - 10/ -8 • -2 = +16\]

OpenStudy (anonymous):

So, \[0 = n^2 - 8n - 2n + 16\]

OpenStudy (anonymous):

I'm alone again...I can't wait to get on the road again! :D

zepdrix (zepdrix):

Mmmm ok if you're going to break it up like that, I guess we can factor by grouping.

OpenStudy (anonymous):

I found my answer! Would you like to see the process?

OpenStudy (anonymous):

\[n^2 - 8n\] \[n^2, n\] \[n ( n - 8 )\]

zepdrix (zepdrix):

\[\Large\rm 0 = \color{royalblue}{n^2 - 8n} \color{orangered}{- 2n + 16}\]\[\Large\rm 0 = \color{royalblue}{n(n - 8)} \color{orangered}{- 2n + 16}\]k looks good so far,

OpenStudy (anonymous):

\[-2n + 16\] \[-2: 2 • 1\] \[16: 8 • 2\] \[GCF: -2 ( n - 8 )\]

OpenStudy (anonymous):

\[( n - 2 ) ( n - 8 )\]

zepdrix (zepdrix):

Cool. so our zeros are what? n=?

OpenStudy (anonymous):

Our zeroes, good sir/ma'am, are +2, and +8.

zepdrix (zepdrix):

Mmm k cool :) What do you think these values might represent? Read this sentence again: The following function represents the profit `P(n)`, in dollars, that a concert promoter makes by selling tickets for `n` dollars each:

OpenStudy (anonymous):

I think that they represent the amount of dollars the promoter earns from a concert per ticket sold?

zepdrix (zepdrix):

That's the general representation of the function and n. That' doesn't answer part A though. So we set our P(n) equal to zero. So these are the specific ticket prices, `n`, which result in ZERO profit. Understand?

OpenStudy (anonymous):

Oh, so they represent the specific prices, +2 and +8, which will result in a zero profit? Is that my answer with the work as well?

zepdrix (zepdrix):

Yes, part A is that reponse + the work we did. 2$ wasn't a high enough price. We didn't charge enough, didn't make any profit. 8$ per ticket was too expensive, not enough people bought tickets or something.

zepdrix (zepdrix):

Ooo we've gotta complete the square >.<

OpenStudy (anonymous):

Ah yes Part B:....how I am going to not miss doing this. XD

zepdrix (zepdrix):

Going back to our original expression:\[\Large\rm P(n)=\color{royalblue}{-250n^2 + 2500n} - 4000\]We want to complete the square on this blue part.

OpenStudy (anonymous):

Ok

zepdrix (zepdrix):

Again, factoring out a -250 gives us:\[\Large\rm P(n)=\color{royalblue}{-250(n^2 - 10n)} - 4000\]

zepdrix (zepdrix):

The part inside the brackets, do you remember how to complete the square?

OpenStudy (anonymous):

I think I'm supposed to distribute it, right?

zepdrix (zepdrix):

You take half of the b term, and square it. That's the value that completes the square. So we have -10 and we need to cut it half and then square it.

OpenStudy (anonymous):

OOOOohhhhhh.....I was doing this earlier, and I never understood what to do next.

OpenStudy (anonymous):

I remember this part, would you mind helping me with this? It's a little difficult with these numbers, hehe. ^^'

zepdrix (zepdrix):

So half of -10 is -5. Then squaring it tells us that positive 25 is the value we want to complete the square, yes?

OpenStudy (anonymous):

\[P (n) = −250 ( n^2 − 10n + 25 ) −4,000\]

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

\[\Large\rm P(n)=-250(n^2 - 10n\color{orangered}{+25}) - 4000\]Ok good. But do you see the mistake we made? In the land of math, we can't just add 25 willy nilly. We have to keep things balanced, yes?

OpenStudy (anonymous):

Correct, we need to subtract 25 from both sides.

OpenStudy (anonymous):

4,000 - 25 = 3,975

zepdrix (zepdrix):

No no this is where we need to be very careful! We didn't add 25 to the right side. We added 25 inside the brackets yes? ( which results in a different number when taken out of the brackets) So we need to subtract 25 in the brackets as well.

zepdrix (zepdrix):

\[\Large\rm P(n)=-250(n^2 - 10n+25-25) - 4000\]

OpenStudy (anonymous):

Wait...then we do what I said above right? Then-wait. I got it. -4000 - 25. I forgot the negative sign on the 4,000. HA Algebra is a killer! XD

zepdrix (zepdrix):

Our goal is to end up with only this blue part in the brackets:\[\Large\rm P(n)=-250(\color{royalblue}{n^2 - 10n+25}-25) - 4000\]Then we'll be able to write it as a perfect square. So we need to pull the -25 out of the brackets somehow.

OpenStudy (anonymous):

Ok continue...

zepdrix (zepdrix):

No when we pull the -25 out of the brackets, we have to multiply it by the -250 right?

zepdrix (zepdrix):

So it turns into a very different number.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Yeah, like over 5,000 I believe.

zepdrix (zepdrix):

\[\Large\rm P(n)=-250(\color{royalblue}{n^2 - 10n+25})+6250 - 4000\]Yah, not quite over 9000 ;) but yes.

OpenStudy (anonymous):

Haha! I got the 9,000 part! XD I understand what to do!

OpenStudy (anonymous):

So, the answer would be:

OpenStudy (anonymous):

\[P( n ) = -250 ( n^2 - 10n + 25 ) 2,250\]

zepdrix (zepdrix):

\[P( n ) = -250 ( n^2 - 10n + 25 ) +2,250\]Ok good. Now we need to write the bracketed portion as a perfect square.

OpenStudy (anonymous):

Could you show me?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Well we could factor by grouping again if you're more comfortable with that.

OpenStudy (anonymous):

Ok, let's do that!

zepdrix (zepdrix):

\[\Large\rm n^2-10n+25\]

zepdrix (zepdrix):

Do it!! >.<

zepdrix (zepdrix):

brb i need some chocolate milk :U

OpenStudy (anonymous):

Ok...it's good for you heart!

OpenStudy (anonymous):

\[n^2 - 5n - 5n + 25\] \[n ( n - 5 )\] \[-5 ( n - 5 )\] \[( n - 5 )^2\]

OpenStudy (anonymous):

Bum bum!

zepdrix (zepdrix):

\[\Large\rm P( n ) = -250 (n-5)^2 +2,250\]Cool.

zepdrix (zepdrix):

What shape does this function make?

OpenStudy (anonymous):

Hold on, let me think...

OpenStudy (anonymous):

A parabola?

zepdrix (zepdrix):

Mmm ok good, and since the leading term is negative, it's opening downward right?

zepdrix (zepdrix):